OpenStudy (anonymous):

Given: \[\Huge x^{[-4+\log_3(x)]}={1\over 27}\] Attempted: \[\Huge x^{[-4+\log_3(x)]}=3^{-3 }\] \[\huge [-4+\log_3(x)]\cdot \log (x)=-3\cdot \log(3)\] Pretty dumb attempt huh? Multiple choices are: A. \[\LARGE x_1=2.4 \quad ,\quad x_2=2\] B. x=3 ,x=27 C. x=-8 , x=2 D. \[\LARGE x=\frac19 \quad ,\quad x=3\] any hint would be great !

OpenStudy (callisto):

Is it B?

OpenStudy (anonymous):

LOL, I'm not trying to see who's smarter. I don't know the answer :$. I was hoping someone to help me do it :(

OpenStudy (callisto):

Oh sorry, I thought you had the answer... Forgive me.. well I did it like this \[x^{-4+log_{3}x} = \frac{1}{27}\]\[\frac{x^{log_{3}x}}{x^4} = \frac{1}{27}\]\[x^{log_{3}x} = \frac{x^4}{27}\] Then, I did it by observation...

OpenStudy (anonymous):

... give me a minute, I'll try something, although I'm sure I'll screw it up :F

OpenStudy (anonymous):

starting from your second step: \[\Huge {x^{\log _{3}x}\over x^4}=\frac{1}{27}\] \[\Huge 3^3\cdot x^{\log _{3}x}=x^4 \] Now I was thinking about: \[\Huge \log_ab=\frac{1}{\log_ba}\] and \[\Huge x^{\log_xa}=a\] to change exponent of X like... \[\Huge x^{\log_3x}=x^{1 \over \log_x3}\] But I don't know how it would be, since there's a fraction. ... :( maybe : \[\Huge x^{(\log_x3)^{-1}}\]something...

OpenStudy (anonymous):

If I could only get rid of that log exponent that would be much easier ...

OpenStudy (callisto):

Hmm.. You're now doing something I haven't learnt...

OpenStudy (anonymous):

that makes no sense anyway... let's assume we have: \[\LARGE 3^3\cdot 3^{-1}=x^4\] \[\LARGE 3^2=x^4\] \[\LARGE x=\pm \sqrt3 \]

OpenStudy (anonymous):

LOL , I'm trying to do something I haven't learnt too, because I don't see other way :B hahah...

OpenStudy (anonymous):

anyway Log rules still hold ... I didn't invent them !

OpenStudy (callisto):

I'm lost :S Perhaps you can sub the values into the equation to get the answer (This is always the ultimate way for MC question :| )

OpenStudy (anonymous):

I want to know how to do it, anyway... Time to call some1 :F .. thanks a lot ;) @Callisto

OpenStudy (callisto):

Sorry that I can't help :(

OpenStudy (anonymous):

ahha.. If you feel bad that you can't help me, I hate to think how should I feel when I know I never helped you :F

OpenStudy (anonymous):

@satellite73 @Ishaan94 ... please help me ...

OpenStudy (anonymous):

wolframa says that answer is B.. :(

OpenStudy (anonymous):

here is my try first step i guess is take the log (base 3) as you did and write \[(-4+\log_3(x))\log_3(x)=-3\] then multiply out and get \[(\log_3(x))^2-4\log_3(x)+3=0\] solve the quadratic get \[(\log_3(x)-1)(\log_3(x)-3)=0\] \[\log_3(x)=1\implies x = 3\] \[\log_3(x)=3\implies x=27\]

OpenStudy (anonymous):

i didn't check the answers but it looks good if this is one of the choices go with that

OpenStudy (callisto):

*Big applause to satellite73*

OpenStudy (anonymous):

(blush)

OpenStudy (anonymous):

clap clap clap... :F

OpenStudy (anonymous):

I'll just write it down ..

OpenStudy (anonymous):

whoaaa... Sooo sick. I knew you'd help me. Thank you everyone. I'm closing it :)

OpenStudy (callisto):

I'm sure I'm a dumb after reading satellite73's solution :( Thanks for posting this question!

OpenStudy (anonymous):

.. me too, never thought about taking logs in base 3 :F ... can't stop laughing LOL. @Callisto thank you for your time. and everyone who at least thought about helping me. ;)

OpenStudy (anonymous):

my solution is....